40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9480     Entry time: Tue Dec 17 02:10:29 2013
Author: Den 
Type: Update 
Category: LSC 
Subject: locking activity 

Koji, Den

Some results and conclusions from tonight:

PRC macroscopic length is detuned. We measured REFL phases in carrier and sideband configurations - they are different by ~45 degrees for both 11 and 55 MHz sidebands. Additional measurement with phase locked lasers is required.

We got stable lock of PRMI+2arms with CARM offset of ~200 pm. We think this is the point when we should transition to 1/sqrt(TR) signals. We plan to rewire LSC model and also test CM servo with 1 arm during the day.

POP ASC OL shape changes when we reduce CARM offset probably due to normalization by sum inside the PD. Servo gets almost useless when PRMI power fluctuates by a factor of few.

SMA cables were made and installed for the REFL165 RF amplifier in lsc rack.

ELOG V3.1.3-